solve for y:
x+8y= 24
how do you do these? and
solve for w:
4(4-t)=1+w-4

Found 2 solutions by nyc_function, Alan3354:
Answer by nyc_function(2741)

To solve for y means to isolate y on one side of the equation.
x + 8y = 24
8y = -x + 24
y = (-x + 24)/8
y = (-1/8)(x) + 3

==========================================
The same rule applies here. We isolate w.

4(4 - t) = 1 + w - 4
16 - 4t = w - 3
16 + 3 - 4t = w
19 - 4t = w

Answer by Alan3354(69443)

x+8y= 24
how do you do these?
Isolate y, get it on one side of = by itself.
x+8y= 24
Subtract x from both sides
8y = 24-x
Divide by 8
y = 3 - x/8
that's it.
---------
and
solve for w:
4(4-t)=1+w-4
16-4t = w-3
Add 3
19-4t = w
w = 19-4t